OpenStudy (anonymous):

Evaluate the limit ; lim x^3/2 - 8 / x-4 x-> 4 Help Please !

OpenStudy (anonymous):

6 b on this package http://www.la-citadelle.com/courses/calculus/review01.pdf help please !

OpenStudy (anonymous):

Explain the change of base logic please :) !

OpenStudy (anonymous):

I know the answer but need explanation the answer is 3 btw

OpenStudy (anonymous):

you have the question wronge im sorry can you please go to this link ! and see i cannot type the question

OpenStudy (anonymous):

6b please ! http://www.la-citadelle.com/courses/calculus/review01.pdf

OpenStudy (anonymous):

if you want the basic logic: you cannot substitute x = 4 in the denominator because the denominator would then be 0, so you must transform the denominator somehow so its ok to plug x = 4

OpenStudy (anonymous):

then....?

OpenStudy (anonymous):

try consider x^(3/2) - 8 as \[(\sqrt{x})^{3}-2^{3}=...\]

OpenStudy (anonymous):

Oooh okay .... mind guiding me all the way please :$ ! <333

OpenStudy (anonymous):

and x-4 = \[=(\sqrt{x})^{2}-2^{2}=...\] I think you can cancel out some terms after that I find your lim

OpenStudy (anonymous):

it's some writing... give me few min

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

here we go: \[=((\sqrt{x})^{3}-2^{3})/((\sqrt{x})^{2}-2^{2})=\]

OpenStudy (anonymous):

\[=(\sqrt{x}-2)*((\sqrt{x})^{2}+2\sqrt{x}+4)/(\sqrt{x}-2)(\sqrt{x}+2)=\] do you see it now?

OpenStudy (anonymous):

ooohh difference of cubes. ..

OpenStudy (anonymous):

\[(x+2\sqrt{x}+4)/(\sqrt{x}+2)=\] now get your x->4 you shouldn't have problems any more (no 0/0) :))

OpenStudy (anonymous):

omg omg omg yeah i get it :$ thanks hunnie ! <3 can you help me with 6c :$ ! <333 xoxoxoxoxox

OpenStudy (anonymous):

yes, lim should be =3 6c?

OpenStudy (anonymous):

from here http://www.la-citadelle.com/courses/calculus/review01.pdf

OpenStudy (anonymous):

I think that exactly (!) the same approach will work! Top: \[=(x^{1/6})^{3} - 1^{3}=(x ^{1/6}-1)(x^{2/6}+x ^{1/6}+1)\]

OpenStudy (anonymous):

Numerator: \[(x ^{1/6})^{2} - 1^{2}=...\] can you finish it?

OpenStudy (anonymous):

yess bozzz :) thankyou <3

OpenStudy (anonymous):

after "unfolding" & canceling out you'll get (3/2) Good Luck!

OpenStudy (anonymous):

want my number babe :) ?

OpenStudy (anonymous):

tehe i jokethanks :)

OpenStudy (anonymous):

i was thinking....too long :( i'll survive :)